Solve the boundary value problem

y''-y=0

With boundary condition

y(0)=0, y(2)=3.6268


1
Expert's answer
2021-05-20T11:45:38-0400

Given,

y''-y=0


"(D^2-1)y=0"


Auxiliary equation is

"m^2-1=0\\(m-1)(m+1)=0



 \\m=1,-1"


Solution of the equation is-

"y=c_1 e^x+c_2 e^{-x}"


As "y(0)=0\Rightarrow 0=c_1+c_2~~~~-(1)"


Also, "y(2)=3.6268\Rightarrow 3.6268=c_1 e^2+c_2e^{-2}"


"7.389c_1+0.1353c_2=3.6268~~~~~~-(2)"


Solving equation (1) and (2) and we get-

"c_1=0.5,c_2=-0.5"


So Solution is-


"y=0.5e^x-0.5e^{-x}"